# Generative Consistency Models for Estimation of Kinetic Parametric Image Posteriors in Total-Body PET

### v1

@@ 4 Discussion
This study describes a computational framework that produces fully Bayesian voxel-wise kinetic parameter maps, including full posterior distributions, across an entire TB-PET volume. By recasting posterior sampling as a conditional consistency mapping problem, the proposed CM reduces the hundreds to thousands of reverse diffusion steps required by DDPM and SBD into a single forward pass followed by a brief, fixed refinement schedule. This idea draws on the theoretical foundations of consistency models in computer vision [17] and extends them to medical imaging inverse problems at a scale that has not been reported previously in PET. This method is a fast and reliable alternative to classical Bayesian tools such as approximate Bayesian computation (ABC) [29] or Markov-chain Monte Carlo (MCMC) sampling [8] whose computational demands have so far confined rigorous posterior estimation to a limited number of regions of interest.
A systematic comparison with the classical approaches shows that CM preserves statistical accuracy but is about five orders of magnitude faster than MCMC. Across 100 synthetic test samples, posterior means for the micro-kinetic parameters deviated by <5<5% from MCMC estimates, whereas ABC produced broader tails and required tens of millions of forward simulations per voxel. Processing 1 slice of TB-PET data (66,154 voxels) with CM completed in 1 min, compared with 2 min for parallel ABC on the same GPU cluster and >1>1 month for parallel MCMC on a 128-core CPU cluster. Additionally, the CM completed processing the entire TB-PET data (4.4 million voxels) within 1.25 hours. Our study thus paves the way for routine Bayesian whole-body parametric imaging with uncertainty estimation.
We also demonstrated that CM outperforms modern deep-learning baseline methods. DDPM and SBD match the prediction accuracy of CM but require hundreds to thousands of reverse diffusion steps, which significantly slows down the sampling. CVAE systematically underestimates variance mainly because the objective function minimizes the forward Kullback–Leibler term, which harshly penalizes allocating probability mass to regions where the true posterior is low [33, 34]. GAN suffers from biased posterior means and underestimated credible intervals, since the generator need only fool the discriminator on the regions it covers; this incomplete coverage whether by missing modes or by over-concentrating on part of a single mode shifts the empirical mean and leads to narrow credible intervals [35, 36].
@@
Two systematic discrepancies emerge when contrasting the Patlak derived and CM derived $K_{i}$ maps. First, Patlak slopes are uniformly attenuated relative to CM estimates because the measured tissue curves are not corrected for vascular spill‑in. Therefore, the resulting Patlak slope corresponds to $(1\!-\!V_{b})\,K_{i}$, so even modest blood‑volume fractions (3–5 % in most tissues, higher in highly perfused organs) translate into lower $K_{i}$ values. Second, the spatial correspondence between Patlak and CM maps is imperfect. Patlak relies on the assumption of irreversible tracer trapping ($k_{4}=0$) to fit a straight line at mid-late time points [43]. This assumption is violated in organs with predominantly reversible kinetics, and elevated noise in low‑count voxels further destabilises the fit, thus introducing artefacts and regional bias [44].
@@
Although the present network was trained solely on FDG data, its modular encoder–decoder architecture makes adaptation to other tracers, such as 18F‐FET for amino‐acid metabolism, 18F‐FMISO for hypoxia imaging and 11C‐raclopride for neuroreceptor studies—straightforward. Parameter efficient techniques such as adapters and low‐rank (LoRA) fine‐tuning [45, 46] can graft tracer‐specific kinetics onto the existing backbone with only a few thousands training samples. This is yet to be demonstrated but will be investigated in the next stage of this research.
Several limitations of our study warrant acknowledgment. First, the training dataset covered a limited range of noise levels based on eight subjects. Acquisition settings with substantially higher noise (e.g., ultra-low-dose protocols, motion corruption) may exceed the model’s learned range, requiring retraining or fine-tuning with synthesized high-noise samples, or prior denoising. Second, our framework does not currently model organ-specific delays, i.e., differences in the arrival time of the tracer between organs within the same patient. Implementing this extension would require organ-specific AIFs, for example derived from imaging data. We acknowledge this as an important direction for future work. Finally, reversibility decisions rely on 95 % credible intervals. More informative visualizations, such as voxel‐wise probability maps for $k_{4}>0$ or clinically weighted thresholds could yield richer insights.
@@ 5 Conclusion
We present PET-TRACER, a consistency model framework that generates voxel-wise kinetic parameter maps, posterior probabilities, and model selection maps from dynamic TB-PET data. The method matches MCMC accuracy, performs competitively with recent deep-learning approaches, and greatly reduces computation time. It is particularly suited for studying kinetic heterogeneity in multi-organ TB-PET and systems physiology. Future work will extend to other tracers and adapt for low-dose or motion-corrupted scans. PET-TRACER is available on GitHub: https://github.com/yundumbledore/PET-TRACER.
@@ Acknowledgment
The authors would like to thank Paul Roach, Dale Bailey, Sally Ayesa and Liz Bailey for their advice on protocol design and the nuclear medicine staff at Royal North Shore Hospital for performing the TB-PET studies. We also acknowledge the facilities and the scientific and technical assistance of Sydney Imaging, a core research facility at The University of Sydney, and the National Imaging Facility, a National Collaborative Research Infrastructure Strategy (NCRIS) Capability.

### v2

@@ 4 Discussion
This study describes a computational framework that produces fully Bayesian voxel-wise kinetic parameter maps, including full posterior distributions, across an entire TB-PET volume. By recasting posterior sampling as a conditional consistency mapping problem, the proposed CM reduces the hundreds to thousands of reverse diffusion steps required by DDPM and SBD into a single forward pass followed by a brief, fixed refinement schedule. This idea draws on the theoretical foundations of consistency models in computer vision [21] and extends them to medical imaging inverse problems at a scale that has not been reported previously in PET. This method is a fast and reliable alternative to classical Bayesian tools such as ABC [33] and MCMC [12] whose computational demands have so far confined rigorous posterior estimation to a limited number of regions of interest. A systematic comparison with the classical approaches shows that CM preserves statistical accuracy but is about five orders of magnitude faster than MCMC. Across 100 synthetic test samples, posterior means for the micro-kinetic parameters deviated by <5<5% from MCMC estimates, whereas ABC produced broader tails and required tens of millions of forward simulations per voxel. Processing 1 slice of TB-PET data (approx. 83,000 voxels) with CM completed in 1 min (using 4 NVIDIA A100 GPUs), compared with 2 min for GPU-accelerated ABC (1 NVIDIA A100 GPU) and over 1 month for parallel MCMC on a 128-core CPU cluster. Additionally, the CM completed processing the entire TB-PET data (4.4 million voxels) within 1.25 hours on the 4-GPU setup. Our study thus paves the way for routine Bayesian whole-body parametric imaging with uncertainty estimation.
We also demonstrated that CM outperforms modern deep-learning baseline methods. DDPM and SBD match the prediction accuracy of CM but require hundreds to thousands of reverse diffusion steps, which significantly slows down the sampling. CVAE systematically underestimates variance mainly because the objective function minimizes the forward Kullback–Leibler term, which harshly penalizes allocating probability mass to regions where the true posterior is low [41, 42]. GAN suffers from biased posterior means and underestimated CrIs, since the generator need only fool the discriminator on the regions it covers; this incomplete coverage whether by missing modes or by over-concentrating on part of a single mode shifts the empirical mean and leads to narrow CrIs [43, 44].
@@
The clinical utility of uncertainty estimation enables more robust decision-making, particularly in the high-noise environments inherent to voxel-level total-body PET. In these noisy conditions, the inverse problem of kinetic modeling becomes ill-posed. The likelihood surface is often flat or complex, causing standard point estimators (e.g., NLS) to exhibit high variance or converge to local minima. In contrast, the CM provides the full posterior distribution, capturing the uncertainty of model parameters due to noisy measurements. We demonstrate the practical value of this via the voxel-wise model selection (Fig. 10), where posterior distributions differentiate reversible from irreversible kinetics. A further plausible clinical translation is the use of posterior variance as a voxel-wise reliability index, enabling clinicians to distinguish high-confidence metabolic features from noise artifacts that require further investigation.
@@
Two systematic discrepancies emerge when contrasting the Patlak derived and CM derived $K_{i}$ maps. First, Patlak slopes are uniformly attenuated relative to CM estimates because the measured tissue curves are not corrected for vascular spill‑in. Therefore, the resulting Patlak slope corresponds to $(1\!-\!V_{b})\,K_{i}$, so even modest blood‑volume fractions (3–5 % in most tissues, higher in highly perfused organs) translate into lower $K_{i}$ values. Second, the spatial correspondence between Patlak and CM maps is imperfect. Patlak relies on the assumption of irreversible tracer trapping ($k_{4}=0$) to fit a straight line at mid-late time points [49]. This assumption is violated in organs with predominantly reversible kinetics, and elevated noise in low‑count voxels further destabilises the fit, thus introducing artefacts and regional bias [38]. While NLS is often considered a standard estimator for the full two tissue compartment models, it is known to exhibit bias, particularly when inequality constraints are applied to model parameters [50]. In contrast, MCMC provides an asymptotically unbiased estimate of the posterior distribution. By training the CM to approximate the MCMC posterior, our method aims to inherit this unbiased property while bypassing the computational cost that renders MCMC impractical for total-body PET.
Many implementations of model selection metrics, such as AIC, typically rely on the residual sum of squares (RSS) derived from fitting competing models [51, 52, 53]. However, using this RSS-based formula implicitly asserts that the noise follows a Gaussian distribution—an assumption that is often inaccurate for reconstructed images. Furthermore, these methods rely on point estimates (MLEs) and asymptotic assumptions that are unstable in high-noise, low-count voxel data [54]. In contrast, our proposed framework enables model selection via a direct probability statement in a single inference pass. Being likelihood-free, it avoids the need for valid noise assumptions or arbitrary complexity penalties. Instead, we inspect the 95% CrI of $k_{4}$. If the CrI for $k_{4}$ includes zero, the data is deemed consistent with irreversible kinetics. This allows for probabilistic model selection that naturally integrates parameter uncertainty.
@@
Although the present network was trained solely on FDG data, its modular encoder–decoder architecture makes adaptation to other tracers (e.g., [18F]FET, [18F]FMISO) or alternative kinetic structures (e.g., one-tissue or dual-input models) straightforward. Parameter efficient techniques such as adapters and low‐rank (LoRA) fine‐tuning [55, 56] can graft tracer‐specific or model-specific kinetics onto the existing backbone with only a few thousands training samples. This is yet to be demonstrated but will be investigated in the next stage of this research.
The selection of the CM was motivated by the specific topological properties of kinetic analysis. Unlike high-resolution image generation where the data manifold is highly curved, often necessitating trajectory straightening techniques like Rectified Flow [57] or Flow Matching [58], voxel-wise 1D kinetic estimation involves a topologically simpler manifold where the direct mapping learned by CM is highly effective. However, we acknowledge that the field is rapidly evolving. Recent innovations such as Improved Mean Flows [59] and Continuous-Time Consistency Models [60] offer enhanced training stability and prediction accuracy. Future work will explore extending the current framework with these advanced methods.
Several limitations of our study warrant acknowledgment. First, the training dataset covered a limited range of noise levels based on eight subjects. Acquisition settings with substantially higher noise (e.g., ultra-low-dose protocols, motion corruption) may exceed the model’s learned range, requiring retraining or fine-tuning with synthesized high-noise samples, or prior denoising. Second, regarding generalizability across different scanners and reconstruction protocols, we note that while the current model is robust to a range of noise magnitude (due to the wide noise range $\lambda\in[3,7]$ used in training), it assumes a specific noise texture and temporal framing (i.e., sampling schedule). A ’universal’ model capable of handling arbitrary framings or reconstruction kernels is beyond the scope of this work. However, the proposed framework serves as a foundational model and adaptation to different scanner geometries or reconstruction-induced noise correlations can be achieved via parameter-efficient fine-tuning (e.g., LoRA adapters [56]). Third, regarding tracer arrival times, our CM framework is inherently capable of handling variable delays, having been trained on offsets ($t_{0}$) ranging from 0 to 1.25 minutes. However, the current clinical application utilizes a single global input function derived from the aorta. Consequently, while the CM is robust to temporal misalignment (as shown in Fig. 5), it does not explicitly model the specific physiological delay and dispersion differences between organs, as this would require organ-specific AIFs. Implementing such an extension by deriving local input functions from the imaging data remains an important direction for future work. Finally, this study applied the two tissue compartment model as a universal kinetic structure, while specific regions (e.g., lungs or bone) may benefit from alternative models. Future work will extend this framework to perform Bayesian model selection together with posterior estimation of kinetic parameters.
@@ 5 Conclusion
We present PET-TRACER, a consistency model framework that generates voxel-wise kinetic parameter maps and model selection maps from dynamic TB-PET data. The method matches MCMC accuracy, performs competitively with recent deep-learning approaches, and greatly reduces computation time. It is particularly suited for studying kinetic heterogeneity in multi-organ TB-PET and systems physiology. PET-TRACER is available on GitHub: https://github.com/yundumbledore/PET-TRACER.
@@ Acknowledgment
The authors thank Paul Roach, Dale Bailey, Sally Ayesa and Liz Bailey for their advice on protocol design and the nuclear medicine staff at Royal North Shore Hospital for performing the TB-PET studies. We also acknowledge the facilities and the scientific and technical assistance of Sydney Imaging, a core research facility at The University of Sydney, and the National Imaging Facility, a National Collaborative Research Infrastructure Strategy (NCRIS) Capability.